Calculate the enthalpy of combustion of methane CH4 + O2 --> CO2 + H2O
1
Expert's answer
2020-04-18T06:59:07-0400

CH4 + 2 O2 --> CO2 + 2 H2O

Standard enthalpy of reaction "\Delta"H0r is the sum of formation enthalpies of the products minus the sum of formation enthalpies of reactants. Using the thermodynamic tables of standard formation enthaplies, one finds that:

"\Delta"H0r = ΔH0f(CO2) + 2*ΔH0f(H2O) - ΔH0f(CH4) - 2*ΔH0f(O2) =

= -393.5 + (-2*285.8) - 2*0 - (-74.9) = -890.2 kJ

Answer: entalpy of combustion is -890.2 kJ